Magnesium chloride commonly occurs as a hexahydrate. a. (5 pt) Write the chemical formula for magnesium chloride hexahydrate. (10 pt) Calculate the formula weight for magnesium chloride hexahydrate. Express your answer in atomic mass units to the nearest hundreth of an amu Show your work. b. 0.01 amu). (10 pt) Calculate the percent by mass of magnesium chloride in magnesium chloride hexahydrate. Express your answer in % as three significant digits. Show your work. c. (10 pt) Calculate the percent by mass of water in magnesium chloride hexahydrate. Express your answer in % as three significant digits. Show your work. d.
Answer #1

Ans :

a) The chemical formula of the compound is given as : MgCl2.6H2O

b) The formula weight of the compound is calculated by adding the molecular mass of all the atoms , so here the formula weight of the compound magnesium chloride hexahydrate is given as :

= 24.305 + 2(35.453) + 6(15.999) + 12(1.008)

= 203.30 amu

c)

Percent by mass of magnesium chloride = ( mass of MgCl2 / mass of MgCl2.6H2O) x 100

= (95.21 / 203.30) x 100

= 46.8 %

d)

Percent of water = 100 - 46.8

= 53.2 %